OpenStudy (anonymous):

help please! in a sample of 26 hand held calculators, 20 are known to be non functional. if 6 of these calculators are selected at random, what is the probability that exactly four in the selection are non functional?

OpenStudy (campbell_st):

looks like binomial probability p = 20/26 and q = 6/26 P(non function = 4) = 6C4 (20/26)^2 x (6/26)^4

OpenStudy (anonymous):

well the options are a. .667, b. .316, c. .769, d. .300, and e. 0 i didn't get any of those when i did the equation